### v1

# Empirical Analysis of Pull Requests for Google Summer of Code

Abstract
Internship and industry-affiliated capstone projects are popular ways to expose students to real world experience and bridge the gap between academic training and industry requirements. However, these two approaches often require active industry collaboration and many students often struggle to find industry placements. Open-source contributions is a crucial alternative to gain real world experience, earn publicly verifiable contribution with real world impact, and learn from experienced open-source contributors. The Google Summer of Code is a global initiative that matches students or new contributors with experienced mentors to work on open-source projects. The goal of the program is to introduce the students to open-source, help gain valuable skills under the guidance of a mentor, and hopefully continue to contribute to open source development; thereby, provide a continuous pool of talented new contributors necessary for maintaining an open source project. This study presents an empirical analysis of pull requests created by interns during the Google Summer of Code program. We extracted and analysed 17,232 pull requests from 2456 interns across 1937 open-source projects. The results show that majority of the tasks involve both code-intensive tasks like adding new features and fixing bugs as well as non-code tasks like updating the documentation and restructuring the code base. The feedback from reviewers covers code functionality and programming logic, testing coverage, error handling, code readability, and adopting best practices. Finally, we discuss the implications of these results for software engineering education.
Index Terms:
open source, pull requests, summer programs, review comments.
I Introduction
Majority of the software engineering education research are usually directed towards students within formal academic setting. Unfortunately, these traditional approaches to software engineering education often limit students’ exposure and engagement to real-world projects; thereby, failing to fully harness their potential and creativity [1],[2]. The industry has often noted that fresh graduates have limited exposure to real-world projects and are often ill-equipped to deal with industrial projects [3],[4],[5],[6]. A number of approaches such as internships and university-industry collaboration on capstone projects have been designed to expose students to real-world projects and provide valuable industrial experience. However, these initiatives often require an active industry partner, and this may not be feasible in all places at all times.
In the same vein, open-source software communities are continuously in need of new contributors to maintain and sustain the code base. Contributions to open source can be publicly verified and may be used to demonstrate competence during a job search. Furthermore, open-source code repositories are abundant, easy to access, and provide real-world examples that are needed to prepare students for the industry. Therefore, recruiting students to work on open-source code repositories will equip students with the necessary industrial experience, provide publicly-verifiable evidence of competence that can be useful for students during a job search, and help with the needed pool of contributors necessary to maintain an open-source code base.
The Google Summer of Code (GSoC) [^1] https://summerofcode.withgoogle.com/[7] is an annual, global three-month internship program that connects students to open-source software communities. The program provides students the opportunity to work on real-world open-source projects, with guidance from experienced mentors in the field. Google provides stipends to student participants after completing certain milestones. The program is the first global initiative to introduce students to open source software development and bring in new contributors. Since 2005, the program has connected more than 19,000 new contributors from 112 countries with 18,000 mentors from 133 countries. The program has generated over 43 million lines of code across numerous projects for more than 800 open-source organizations [^2] https://summerofcode.withgoogle.com/about. The code contributions of the participants are publicly available. Therefore, the GSoC program can provide valuable data to gather insights into activities of novices in open-source development.
Contributions to open-source are often done via pull requests where a contributor first clone the code base, add changes (e.g, add a new feature) to the local copy, and then notifies the repository team that they have completed a feature or modified the code base [8]. Collaborators can then review and provide comments on the proposed changes. Code review comments provide valuable feedback that guide improvements in the software development process. A maintainer has to approve the requested change before they can be integrated into the main code base. Although a number of studies have been conducted on pull requests, there is limited research on the analysis of pull requests by interns or students such as pull requests of participants in the Google Summer of Code program.
This study provide an analysis of pull requests created by interns during the Google Summer of Code program from 2020 to 2023 (a 4-year period). We analyze pull request data, review comments, project-related features and contributor (intern) data to provide informative insights into the kind of contributions the intern’s make, the nature of feedback they receive, and possible implications for software engineering education. The study analyzed 17232 pull requests by 2456 interns across 1937 projects. We then used topic modelling techniques to gain insight into the common themes in the pull request title, description and review comments. In summary, the paper provides the following contributions
  1. 1.
    A publicly available dataset  [^3] https://github.com/compedutech/gsoc of 17,232 pull requests created by GSOC interns from 2000 to 2023. These pull requests were extracted directly from Github via the GiHub API. The scripts used for web scrapping (to identify the GitHub URL) and the script for extracting the pull requests from GitHub are also provided.
  2. 2.
    An insight into the success or failure of the GSOC program in promoting open-source contribution. Specifically, we provide information on further contributions by the interns to the same project after the GSOC program.
  3. 3.
    An analysis into the type of task the interns worked on and the feedback they receive from the other project collaborators.
  4. 4.
    A discussion on the possible implication of the study results for software engineering education.
II Background on Google Summer of Code Program
Google Summer of Code (GSoC) is an annual, global program run by Google that encourages university students or new contributors to participate in open-source software development. The program pairs new contributors with mentors from open-source organizations. Google provide stipends to contributors who successfully meet some specified milestones. The following paragraph provides more detailed overview of the program.
  • Goal. GSoC is designed to introduce students to open-source software development by working on real-world projects with guidance from experienced mentors. The program has two main goals. Firstly, to expose students or novice to open-source development under the guidance of experience mentors and gain real-world experience. Secondly, to provide open-source organizations with new contributors that can continue to contribute to the open source communities long after the program is completed.
  • Eligibility. Participants must be students or new to open-source from all over the world. They must be atleast 18 years old, eligible to work in their country of residence, and residence country should not be currently embargoed by the United States.
  • Organization. Open-source organizations apply to be mentoring organizations, offering projects for students to work on. The organization must have an active open source project and not be based in a currently embargoed by the United States.
  • Mentorship. Experienced members of open-source organizations mentor students throughout the program.
  • Timeframe. The standard timeline for the program is 12 weeks and it usually runs over the summer ( from May to August). However, the mentors and contributors can agree to extend the program till 22 weeks.
  • Application process. First, open-source organizations submit proposals for specific software development projects. Students would then create proposals explaining how they would approach the project. The mentoring organizations would then select students based on their proposals and the skills demonstrated.
  • Onboarding and mentorship. Selected students are paired with mentors from the open-source organization. The student spend the first few weeks to get familiar with the code base and work on the expected milestones in collaboration with their mentors. The students then spend the next twelve (or more) weeks working on the implementation of the proposed solution and completion of the milestone.
  • Completion. Students work on their projects throughout the summer, delivering code, features, or bug fixes, and submitting reports to both their mentors and Google.
  • Stipend. Google provides stipends to students based on their location and the duration of the program. The payment of the stipend are often contingent on the completion of some specified milestones.
The GSoC program provides an excellent opportunity for students to develop their skills and make meaningful contributions to projects used by people around the world. There are numerous benefits associated with the program and they include the following [9],[10].
  1. 1.
    Learning opportunity. Students gain hands-on experience in software development in a real-world context.
  2. 2.
    Networking. Participants build relationships with mentors and other open-source contributors.
  3. 3.
    Career development. The contributions implemented during the GSoC program are public and easy to verify. Many students saw the GSoC program as an opportunity to build their resume, kickstart their career, and increase their visibility when seeking a job [9]. Furthermore, the program is popular, highly regarded by employers, and many students go on to work for the organizations they contributed to. This makes participation in the program an excellent component of the participants resume when searching for jobs.
  4. 4.
    Open-source contribution. Participants contribute code to open-source projects, which is publicly available and can be used by the global software community.
  5. 5.
    Open-source sustainability. The program provides new contributors to open-source communities which is necessary for maintaining the open-source projects. Tan et al. [10] noted that sustainability of the project is a major motivating factor for mentors to participate in the program.
The GSoC program offers an excellent avenue to introduce students or newcomers to open source, and support the sustainability of open-source communities. Majority of the organizations participating in the program have also reported that the program helps them to find new contributors who are active and committed to the open source project [11]. The contributions of the students to the open source community are also public and accessible to anyone. This public data provides valuable dataset for analyzing the impact of the program, understand the challenges or barriers to open source software development, and probe factors that ensure successful engagement with the open source communities.
III Literature Review
A number of research has been conducted on summer code programs and pull requests, but to the best of our knowledge, none has provided empirical analysis of the pull requests developed during these summer programs. Silva et al. [9][12] conducted a qualitative study to understand what motivated students to participate in the GSoC program and then developed a theoretical framework to model student engagement. Their findings show that students often participate in the program to gain new software engineering skills, enhance their resume, and earn stipends. Tan et al. [10] also conducted a qualitative study on mentor’s participation in the GSoC program. The results of their analysis shows that intristic motivation to sustain the open source code base, gain new skills, and increased recognition in the open-source community are the major reasons why mentors participate in the program. Trainer et al. [13] conducted a study of 22 GSoC projects and 22 hackatons to understand student engagements with the open-source scientific community. Their findings reveal that private communications with the mentor often create strong ties with the mentor while public sharing of progress report creates weak ties with other community members. They also discovered that about 18% of the students later became mentors for others.
Majority of the approach so far have used qualitative study (without considering the code base) to analyze students and mentors participation in the GSoC program. The closest to our work was conducted by Silva et al. [14][15] who analyzed the code commits of 866 students in the GSoC program from 2013 to 2015. The results show that their is a strong correlation between the number of commits a student produce and the retention of that student to the project during and after the GSoC program. They also noted that 82% of the projects merged atleast one commit of the students to their code base. Commits provide low-level data on the students contribution as they provide information on how and why each line of code was changed. Hence, it may be challenging to analyse the kind of tasks or features they work on. Pull requests on the other hand, contains aggregate changes from multiple commits and provide a high-level data on the kind of tasks the students work on. Pull request also provide opportunities for other members of the community to review the changes, provide comments or feedback, and suggest modifications to the proposed changes [8]. Furthermore, pull requests have been shown to be more precise than commits in evaluating the contributions of community members [16]. Our study extend the work by Silva et al. [14][15] by using pull requests to understand student’s contribution to the GSoC program.
There has also been studies conducted on pull requests. Gousios et al. [8] conducted an exploratory study of the pull based model on 291 projects. They found out that the pull based model reduces the time needed to merge contributions to the code base, increases the development turnover, and provided more opportunities for community members to engage each other. Kalliamvakou et al. [17] also conducted a study on GiHub repositories and found out that pull requests are often used for code reviews and issue resolution. liu et al [18] also noted that Pull request can lead to more active code repositories. Yu et al. [19] and Kerzazi et al. [20] developed theoretical frameworks for suggesting reviewers to review pull requests. Zhang et al. [21] examined competing pull requests and noted that 31% of pull requests in 75% of the repositories they studied tend to modify the same code during the same period. Jiang et al. [22] studied reopened pull requests and noted that reopened pull requests often had more comments and lower rate of being merged with the code base than pull requests that have not been reopened. Legay et al [23] examined the impact of the acceptance or rejection of a pull request on future contribution to the project. Their results show that contributors with higher acceptance rate are more likely to continue contributing to the project while those with rejections are less likely to contribute in the future. The studies discussed in this paragraph have examined pull requests in different context. However, none of the studies have focused on pull requests produced by novice programmers or students. This study extends the literature to provide an in-depth analysis of pull requests from students or new contributors in the GSoC program.
Internships and industry-based capstone projects are a popular way of providing students with real world experience. There has been many studies conducted within the context of internships and capstone projects. Kapoor et al. [24] and Lehman et al. [25] reported that internships often lead to improved professional development, while Groeneveld et al [26] noted that both internships and capstone projects promotes the development of soft skills among students. Dean et al . [27] noted that the impact of internships on students depend on the role of the student during project execution. Sudol et al. [28] also reported that students who had completed at least one internships tend to have reduced misconceptions about software engineering concepts. The related work on internships discussed so far have adopted qualitative studies to analyze the impact of internships on students without looking at the actual code contributions. The closest work to our study was by Menezes et al. [29] where they described their experience on an internship model that pairs students with volunteer mentors in open source projects. The authors then conducted a qualitative study using surveys and reports to understand the impact of the internship model on the student skills and job placement. Their results show that significant increase in student’s technical skills, soft skills, and job placement after graduation. They also noted that students who are paid stipends are more willing to work full time on the project.
This literature review section have presented a number of works related to GSoC, pull request and internships. However, none of the work examined has conducted a large-scale study on pull requests within the context of internships or summer coding programs. This study extends the literature review by analyzing the pull requests of students or interns participating in the GSoC program.
IV Design Methodology
This section discusses the data collection process, the research methodology adopted in this paper, and the research questions we aim to answer.
IV-A Data Collection
We extracted pull request data via web scrapping and the GitHub API. The following paragraphs highlights how we extracted the needed data.
  1. 1.
    Firstly, we used the Selenium Python Package to scrape the URLs of each intern’s personal reflection website from the GSOC web page for each year (2020, 2021, 2022 and 2023).
  2. 2.
    Secondly, we extracted all the URL links from each intern’s webpage (by searching for ”a” html tag).
  3. 3.
    Thirdly, we filtered out the links containing ”github” and (”issues” or ”pull”). For example the url link ”https://github.com/repos/asyncapi/community/pulls/720” contains the words ”github” and ”pull”. Most pull requests url on github often follow this pattern and we believe is the easiest way to extract most of the pull request link.
  4. 4.
    Finally, we used the GitHub API to extract the pull request data associated with the filtered links.
We extracted three main types of data related to the pull request, the contributor (intern) that created the pull request, and the project (or code repository) associated with the pull request. At the end of the Data collection process, we collected 17,232 pull requests from 2080 interns across 1937 open-source projects. The following subsections provides statistical discussion of the pull requests, projects and users embedded in the data.
IV-A1 Pull Requests Data
The pull request data has 22 features related to different aspects of the pull requests.The features include the title (title of the pull request), body (description of the pull request), is_merged (if the pull request has been merged to the code base), created_at, closed_at, review_comments, commits, and so on. We also added a derived features that were computed from the the original data. The derived features is time_taken which indicate the time difference between when the pull request was open and closed.
The final pull request data shows that there were a total of 17232 pull requests with 60 million lines of code added, 18 million lines of code deleted, across more than 300,000 files. On the average, a pull request was opened for 4.7 hours, has 4.3 comments and 9.5 review comments. Table I provides a statistical description of key pull request features. The is_merge feature is assigned a value of 1 if the pull request has been merged and a value of 0 otherwise. Similarly, the State feature is assigned a value of 1 if it is closed, else it is assigned a value of 0. The table shows that 14775 (or 85.7% of 17232) pull requests were successfully merged with the code base while 16235 (94.7%) pull requests have been resolved and currently closed.
Statistics Is_merged State Additions Deletions Changed Files Commits PR comments Review Comments Time taken (Hours)
Total 14775 16325 60,049,227 18,668,966 353,946 198,880 74, 343 165,330 14,898,582
Mean 0.857 0.947 3484.95 1083.45 20.5 11.5 4.3 9.5 4.7
Std Dev 0.35 0.22 67352.30 52568.90 263.91 84.04 8.48 26.08 2812
TABLE I: Descriptive Statistics of Pull Request
Statistics Forks Stars Watchers Contributors Commits Issues Branches Pull Requests
Total 1104727 3479159 117861 335513 14141978 328060 66728 53210
Mean 570.62 1797.09 60.88 173.30 7304.74 169.45 34.47 27.48
Std Dev 3132.37 8642.19 265.01 794.45 34255.74 864.12 142.84 131.92
TABLE II: Descriptive Statistics of Projects
Variables Number of Pull Accepted Pull Reject Pull Projects Public Repositories Following Followers
Total 17231 14775 2456 3117 91957 66295 124813
Mean 8.28 7.10 1.18 1.50 45.32 32.67 61.51
Std Dev 10.37 9.72 2.83 1.54 47.30 76.84 287.95
TABLE III: Descriptive Statistics of Users
IV-A2 Projects Data
The project data has 8 features including the project name, programming language. The other features are the number of Forks, Stars, watchers, contributors, commits, branches, issues, and pull requests. The total number of projects extracted from GSoC data was 1937.
Table II provides a statistical overview of the project metadata. The table shows that the projects the students work on are popular and complex with an average of 1795 stars and have 34 branches on average. The top ten projects based on the number of stars [30] are tensorflow/tensorflow, huggingface/transformers, nodejs/node, rust-lang/rust, godotengine/godot, django/django, opencv/opencv ,tensorflow/models, swiftlang/swift, and webpack/webpack. Furthermore, we noticed that majority of projects was based on the Python programming language. Figure 1 provides a graphical overview of the number of projects associated with the top 20 languages in the dataset.
Figure 1: Number of Projects associated with each Programming Language
IV-A3 Users Data
The user-related data captures information related to each intern that participated in the GSoC program. The user dataset captures 9 features including the contributor id, the number of projects they worked on during the GSoC program, the number of pull requests created during the GSoC program, number of accepted pull requests, number of rejected pull request, and all the review comments associated with the user. Other features include the number of public repositories, followers and followings.
There were a total of 2456 interns in the dataset, including 154 duplicates. The duplicates captures interns who participated in multiple sessions of the GSoC program (e.g., a student who participated in the 2020 and 2021 programs). For this study, the duplicates user are treated as new users because the user contributions (e.g number of pull requests created) will differ from year to year. Table III provides a statistical overview of the user-related data. The number of projects in the table represent the aggregate number of projects the interns worked on and not the unique number of projects. Table II captures the information for each unique project.
IV-B Research Questions
The goal of this study is to provide insights into the pull request activities created by interns during the GSoC program. A detailed analysis of the pull requests can provide insights into the patterns of contributions interns provide in open source repository. We may also be able to understand how collective thinking appears among a group of individuals. We may be able to know the core features the interns often work on, the type of feedback they often receive, and factors that contribute to a successful or unsuccessful GSoC experience. These insights may inform necessary changes to the software engineering curriculum to encourage students to contribute to open source and better prepare students for the workforce. Therefore, the following three research questions guide the study.
  • RQ1. How successful has the GSoC program been in encouraging contribution to open source communities?. A continuous stream of new contributors is crucial to the sustainability of the open source projects. One of the major goals of the GSoC program is to encourage contribution to open-source. This research question aims to unravel how successful the GSoC program has been in attracting and retaining new contributors.
  • RQ2. What are the major tasks that interns often work on? This research questions aims to discover the categories of tasks or contributions the interns often work on. This can provide into the types of tasks interns are mostly interested in, thereby providing a basis for attracting new comers to the open source community.
  • RQ3. What are the major feedback interns often receive on their contributions. The pull request contributions are usually reviewed by experienced contributors or moderators who provide comments on how to improve the contributions of the interns. These feedback can provide valuable insights into the challenges interns face, identify patterns of mistakes or errors by interns, and generate improvements to the software engineering curriculum to better prepare students for the work force and open-source contribution.
These three research questions have been selected to better understand intern’s contribution to the open source, provide guidelines for open source communities on retaining new contributors, and discuss the implications of the findings for software engineering academic curriculum.
IV-C Study Design
For the study, we used topic modelling techniques to answer the research questions. Topic modeling, or topic clustering, is a method that uses unsupervised machine learning techniques to discover topics within a text without human intervention. It is commonly employed to reveal the underlying thematic structures in textual content. This process generally consists of three key steps: pre-processing, topic modeling, and interpretation of results.
Firstly, we pre-process the text documents to reduce improve the accuracy of topic modeling. We converted the text to lowercase and removed irrelevant characters (such as ’/’), punctuations, numbers, and short words with only one or two letters. Then, we tokenize the texts to break the text into individual words, removed stop words (e.g., ’a’, ’an’ ’the’) and stemming to reduce words to their base form (e.g., converting ’understanding’ to ’understand’).
Secondly, we used the Latent Dirichlet Allocation (LDA) [31] from the Gensim Python library [32] to identify topics within the document. We compared coherence scores across a range of topic numbers in order to select the optimal number of topics for each tasks.
Finally, we interpreted the results by associating the generated topics with their respective keywords. We selected posts that contains the identified keywords, manually read the posts to understand the context, and then applied open-coding techniques to label the topics generated from the Gensim package. We adopted a 2-step open-coding process where we first extract the top 20 keywords for each topic identified by the LDA algorithm and then manually assigning labels to the topics based on these keywords. The goal of this final interpretation process is to ensure that we generate a good label that captures the main themes for each topic.
To answer the first research question, we extracted all the pull requests created by the user in the project. We then exclude the pull request that were created before or during the GSoC program. This allows us to examine the contributions of the user to the project before the start of the program and after the GSoC program. It should be noted that we were able to access only the pull request created by the user in the project(s) they worked on during the GSoC program and not other projects. We were unable to track if the user contributed to other open-source projects outside of the project they worked on during the GSoC program.
To answer the second research questions, we applied topic modelling techniques to the title and body of the pull request. The titles conveys the purpose and scope of the changes while the body contains the descriptions of the pull requests. The body of a pull request is crucial for providing context and explaining the changes made in the code, allowing reviewers to understand the scope and rationale behind the work. A well-crafted pull request body typically includes a clear explanation of what the pull request does, the reason for the change, and any special instructions for testing or deployment. Therefore, a detailed analysis of the Pull request title and body (or description) allows us to understand the major tasks or topics the interns work on.
To answer the third research question, we applied topic modelling to the review comments. The review comments in the dataset contains valuable feedback provided on the various pull requests. These comments are essential for improving the code quality, ensuring adherence to best practices, and enhancing clarity and maintainability. A well-constructed review comment provides constructive criticism, suggests improvements, and sometimes includes code snippets to guide developers. An analysis of these comments can unravel the major areas interns need to improve on.
V Results
This section present the results of our analysis with respect to the three research questions discussed in Section IV-B.
V-A RQ1: GSoC Success
The first research questions examines how successful the GSoC program is in promoting open-source contribution. The results of the analysis from Table I show that 85.74% (or 14775 out of 17232) of the pull requests were accepted and successfully merged to the codebase. Furthermore, 1113 out of 2456 users had all their pull request accepted while 319 had no accepted pulls.
Next, we examine the pull requests that were created by each user before and after the GSoC program. The results show that 63.5% (1561 users) contributed at least one pull request to the same project after the program, while 70.7% (1738 users) contributed at least one pull request to the project before starting the GSoC program. This shows that majority of the users started contributing to the project before the GSoC program and many continue to do so after the program. Infact, just 15% contributed only during the GSoC program with no contributions before or after the program. This may indicate that the GSoC program motivates students to contribute to open-source. Students may also have contributed to an open source project before the program to have a higher chance of being selected for the program.
V-B RQ2: Tasks and Contributions
We used word frequencies and topic modelling techniques to analyze the title and description (body) of the pull request with the goal of extracting the type of task the student worked on. First, we extracted the most frequent words that appear in both the title and the description. Figure 2 and Figure 3 illustrates the most recurring word in the title and description.
Figure 2: Frequencies of Words in the Title
Figure 3: Frequencies of Words in the Body (Description)
Next, we used LDA topic modelling to extract the major topics in the description and the key words that capture the topics. Thirdly, we extracted sample text from the documents containing the keywords from the word frequency and topic modelling in order to understand the context of the words. Finally we categorize the topics based on the key words and the context in which they appear. The extraced topics for the tasks are.
  1. 1.
    Adding new features and APIs. This topic involves developing and implementing new features like knowledge panels or API end points. Sample pull request with this topic include: ”Added support for opening a file”, ”Added custom IDE and improved other IDE launch commands”, and ”Add an explore mode to run the app without marking messages as read”.
  2. 2.
    Corrective maintenance tasks. This topic covers pull request that tries to correct an error or ensure a feature is working as expected. Examples of pull request in this topic include ”Fix double-checked locking in ConnectionFactory”, ”Fix axeDev tools + lighthouse error for learner dashboard and make the lighthouse score to 1”, and ”Add and update tests”.
  3. 3.
    Project setup. This topic covers pull requests that deal with initializing the project structure and adding necessary dependencies. Examples of pull request in this topic include: ”Setting up the foundational infrastructure for FastAPI”, ”Code relocated under maryam/”, ”added setup.py, src: setup: Add root install support”, ”Initial PlanetGenerator Setup”, and ”Created the spring boot project with the necessary dependencies”.
  4. 4.
    Documentation update This topic covers pull requests that relate to making changes to the README file or other documentation files to ensure clarity and up-to-date information. Sample pull request are ”Updating or adding information to ReadMe files”, ”Privacy Statement Updates September 2022”, ”Added basic chondro documentation”, ”restructured install documentation”.
  5. 5.
    Refactoring. This topic covers pull requests that aim to change parts of the code base without changing the behaviour. Sample topic include ”wxGUI refactoring: New WorkspaceManager class”, ”refactor: color blending function”, ”refactoring of webapp to consume typesript client”, and ”[GSoC] API change for the newly added system of ODEs solvers”.
The analysis of the second research question show that while majority of the contributions are code intensive, there was a significant part of the pull request that involves changes to the documentation files and restructuring the organization of the code base. This may also indicate the need to emphasize the variety of contributions student can make to open source.
V-C RQ3: Feedback and Review Comments
To answer the third research question, we also applied word frequencies and topic modelling techniques to the review comments to understand the type of feedback reviewers provide for the pull request. The topics identified in the review comments include the following.
  1. 1.
    Code organization and structure improvements. This topic emphasize project structure and organization, particularly where code files, tests, and other assets should be placed. This helps in maintaining a clean project hierarchy, improve readability, and simplify navigation for future developers. Sample comments include ”You should put test in a tests/ directory” and ”If you move this to the test directory, it becomes clearer”.
  2. 2.
    Best practices and coding style. This topic includes comments that suggest changes that align with best practices, such as ensuring minimal exposure of unnecessary components and adhering to consistent coding styles. This helps ensure code reliability, consistency, and security. Sample comments include ”I always prefer if we keep safe and only publish necessary things” and ”Why not just use a str for facet_value?”.
  3. 3.
    Functionality and logic corrections. This topic relates to the clarity of logic or ensuring the code behaves as intended. Reviewers help identify sections where the functionality may not align with the expected outcome or where the logic could be improved. Sample comments include ”This is more the why of knowledge panels” and ”The logic here could be improved by changing this function”
  4. 4.
    Testing coverage and validation. This relates to adding more tests or moving test cases to appropriate locations to ensure the functionality is well-covered. They emphasize the importance of testing edge cases and ensuring that new features are adequately validated before being merged. Sample comments include ”You should write a test for this specific scenario” and ”This becomes easier to test if you move it into a dedicated test case”
  5. 5.
    Error handling and resilience. This topic focuses on ensuring that code is resilient to errors and can handle potential failure cases gracefully. Reviewers point out where additional checks or error handling mechanisms can be implemented to improve robustness. Sample comments are ”Make sure we handle this case properly when the data is missing” and ”Consider adding a check for this specific error”.
  6. 6.
    Code readability and clarity. This topic relates to request for improvements in code readability, suggesting changes that make the code more understandable. This involves better naming conventions, simpler logic flows, and clearer documentation or comments. Sample review comments include ”This part of the code could be written more clearly” and ”The variable names should better reflect their purpose”
The result presented above shows some key areas that reviewers have focused on. It can be observed that reviewers are not only concerned about the correctness of the code, but they are also interested ensuring that the newly added or modified code is easy to maintain.
VI Discussion and Implications for Software Engineering Education
The results of the analysis on GitHub pull requests and review comments can have significant implications for software engineering education. The tasks and feedback topics identified in review comments can aid educators to better prepare students for real-world software development practices. The findings highlight key areas that should be emphasized in software engineering curricula, as outlined below.
  1. 1.
    Importance of code reviews in collaborative software development. The study reveals the importance of peer reviews in ensuring the quality of code contributions. For software engineering education, this means integrating code reviews into course projects and emphasizing communication skills.
  2. 2.
    Code organization and structure. The findings reveal that reviewers frequently comment on the structure and organization of code, such as where files should be placed and how modules should be organized. This highlights the need to emphasize code organization in software engineering education by 1) teaching modular design and file organization, and 2) following real-world practices in project layout.
  3. 3.
    Testing and test-driven development The study shows that a significant portion of review feedback is related to missing or inadequate testing. This suggests that educators should place a stronger emphasis on teaching comprehensive testing practices and emphasizing edge cases.
  4. 4.
    Error handling and robustness. Feedback in the review comments frequently addresses issues with error handling, suggesting that students often underestimate its importance. In response, educators should teach best practices for error handling and include error handling in assessment criteria.
  5. 5.
    Performance optimization. Reviewers also suggest performance optimizations or improvement on the coding logic. For software engineering education, this implies teaching algorithmic efficiency and introducing profiling tools.
  6. 6.
    Refactoring and continuous improvement. Reviewers frequently suggest refactoring code for better readability, efficiency, and maintainability. In education, this highlights the need to encourage refactoring as a habit and teach clean code principles.
  7. 7.
    Prepare Students for open-source contributions. Many of the findings reflect practices commonly seen in open-source development environments, such as GitHub. The alignment of software engineering education with these practices can prepare students better to contribute to open-source projects. Educators can use real-world GitHub projects as learning tools, and incorporate the use of version control for collaboration in software engineering classes.
  8. 8.
    Learning to give and receive feedback. The study shows that review comments often provide specific, constructive feedback on how to improve code. Software engineering education should emphasize 1) ability of students to provide constructive feedback and 2) receiving and acting on feedback provided by peers.
The findings of this study offer several key insights for software engineering education. When educators focus on real-world practices with respect to code organization, testing, error handling, performance optimization, and collaboration; they can prepare students better for the challenges they will face in the software industry. The integration of code reviews, refactoring exercises, and open-source contributions into the curriculum will likely foster a culture of continuous improvement and collaboration; thereby, equipping future software engineers with the tools they need for success. These educational adjustments will help students internalize best practices in software development, make them better collaborators, more thoughtful developers, and ultimately more successful engineers in both open-source and proprietary software environments.
VII Threats to Validity
There are several potential threats to the validity of our findings. These threats can affect the reliability and generalizability of the findings. The following subsections discuss the potential threats to validity of the results in terms of construct validity, internal validity, external validity, and conclusion validity.
VII-A Construct Validity
Construct validity refers to whether the measures used in the study accurately capture the concepts or phenomena being investigated. In this case, the key constructs are the themes extracted from pull request titles, descriptions, and review comments. The following represent the construct validity threats.
  • NLP limitations in identifying themes. The natural language processing (NLP) techniques used in the study, such as keyword extraction, tokenization, and topic modelling may not fully capture the nuances and context of review comments. For instance, comments that use implicit language, technical jargon, or sarcasm might be misinterpreted by simple keyword-based approaches. To minimize this threat, we manually validate a sample of comments. However, improving the NLP model by incorporating more advanced models (e.g., transformer-based models) could improve accuracy in capturing nuanced feedback.
  • Pull request descriptions may not fully capture the nature of changes. Pull request descriptions vary widely in quality and completeness. Some pull request may lack detailed explanations of the changes, leading to incomplete or inaccurate summary of tasks.
VII-B Internal Validity
Internal validity refers to the extent to which the observed results can be attributed to the research methods used, rather than external factors. This includes issues that may arise from the dataset or the methods of analysis.
  • Bias in data collection. The dataset used for the study was sourced from only GitHub repositories of projects that participated in the Google Summer of Code program. Since this program in being run by only Google, there is possibility that the projects follow specific coding practices, use particular technologies, or operate in niche domains. Hence, it is possible that the findings may not be representative of broader software development practices.
  • Incomplete review comments. Reviewers may not comment on all issues in the code due to time constraints or priorities. As a result, important problems or suggestions for improvements may be overlooked in the analysis.
VII-C External Validity
External validity pertains to the generalizability of the findings. This involves assessing whether the results of the study can be applied to other software projects, development teams, or domains.
  • Generalizability across different projects and domains. The study may not generalize well to all software projects. Different software projects (e.g., open-source vs. proprietary, small vs. large projects, front-end vs. back-end development) have distinct workflows, contributor behaviors, and review cultures. Therefore, findings from a single type of project may not apply broadly.
  • Varying review practices across teams. Review practices may differ based on team norms, the development process (e.g., agile vs. waterfall), or company policies. A highly structured review process may yield different types of feedback than a more informal review process.
  • Changes over time. The feedback practices and development workflows on GitHub may evolve over time, influenced by the introduction of new tools (e.g., code linters, continuous integration pipelines) or changes in the development community’s culture. This temporal aspect might limit the applicability of findings from historical PR data to current or future practices.
VII-D Conclusion Validity
Conclusion validity concerns whether the conclusions drawn from the analysis are justified by the data and methodology used.
  • Over-interpretation of statistical results. While keyword frequency and thematic analysis offer valuable insights, there is a risk of over-interpreting correlations as causal relationships. For example, a high frequency of comments mentioning ”testing” might not necessarily imply that testing is a major issue across all pull request, but could reflect a focus on testing in certain projects.
  • Unaccounted confounding factors. Confounding factors such as the complexity of the code changes, the experience level of the contributor, or the urgency of the PR could influence the nature of the feedback provided. If these factors are not accounted for, the conclusions may be skewed.
The study provides valuable insights into pull requests and code review comments. However, the validity of the findings could be affected by the aforementioned threats. We have used a variety of approaches (such as manual validation of some of the comments) to minimize these threats and ensure careful interpretation of the results. In the future we will expand the dataset, refine the analytical techniques, and account for variability in development practices to enhance the validity and generalizability of the findings.
VIII Conclusion
This study provides an in-depth analysis of GitHub pull requests and review comments created during the Google Summer of Code program from 2020 to 2023. The study offers insights into common tasks and feedback themes. The results shows that pull requests often focus on feature additions, bug fixes, and documentation updates. The review comments frequently emphasize the need for improved code organization, additional tests, performance optimizations, and better error handling.
These findings can help developers and teams enhance their review practices by focusing on key areas such as test coverage, performance improvements, and maintaining a well-structured codebase. These results also show that there is need to enhance the software engineering curriculum to include emphasis on code reviews, code organization, testing, and best practices from open source repositories in the curriculum.
In the future ,we plan to expand the dataset to include the available dataset for all the years in the GSoC program. We also plan to study the evolution of the tasks and comments across these years. Finally, we plan to explore other summer programs and hackatons that target students to validate the generalizability of this study findings across different repositories and software domains.